OpenStudy (anonymous):

Find the first 5 terms of the sequence a1 = -7 an = an-1+6

OpenStudy (anonymous):

\[a_1=-7\]\[a_n=a_{n-1}+6\] a_2=-7+6=-1 a_3=-1+6=5 a_4=5+6=11 a_5=11+6=17 (By the way, this is called arithmetic sequence with first term -7 and common difference 6)

OpenStudy (anonymous):

Thanks ! for this one: a1 = 6 an+1=-2an+3 is the second term 3/2?

OpenStudy (anonymous):

\[a_1=6\] \[a_{n+1}=-2a_n+3\] \[n=1 -> a_2=-2a_1+3\] \[a_2=-12+3=-9\]